OpenStudy (anonymous):

HEELPP!! The country fair charges $7.00 for adults and $2.00 for children. If $6,000 dollars were collected in one day and there were a total of 2,000 patrons, how many adults and children attended?

OpenStudy (anonymous):

400 adults, 1600 children.

OpenStudy (anonymous):

What are the steps to get that answer though?

OpenStudy (anonymous):

let # of adults attended = x and # of children attended = y 7x + 2y = 6000 x+y = 2000 7x + 2y=6000 -2x - 2y=-4000 (times the equation by -2) 5x=2000 x=400 (# of adults attended) x+y = 2000 y=2000-400 y=1600 (#of children attended)